OpenStudy (anonymous):

How many solutions will the following system of linear equations have? –2x + 4y = –7 A. 0 B. 1 C. 2 D. infinite

OpenStudy (texaschic101):

is there another equation ?

OpenStudy (anonymous):

no

OpenStudy (texaschic101):

then that is not a system

OpenStudy (anonymous):

Just kidding my page didnt load all the way

OpenStudy (texaschic101):

lol...you made me think I was going crazy

OpenStudy (anonymous):

the other equation is y= -1/2x+ 5

OpenStudy (texaschic101):

y = -1/2x + 5 --- slope is -1/2 -2x + 4y = -7 4y = 2x - 7 y = 2/4x - 7/4 y = 1/2x - 7/4 --- slope is 1/2 different slopes, different y intercepts.....means 1 solution just so you know.. same slope, different y intercepts = no solution or 0 same slope, same y intercept = infinite solutions different slopes, different y intercepts = 1 solution